OpenStudy (anonymous):

a kite and its tail total 36ft in length. the tail is five times the length of the body. how long is the kite's tail?

OpenStudy (anonymous):

a kite and its tail total 36ft in length. the tail is five times the length of the body. lets make the kite = k and the tail = t k + t = 36 t = 5k lets plug in ^ for t. k + (5k) = 36 6k = 36 k = 6 ok now we have the length of the kite which is 6 but it is looking for the tail. All we need to do is subtract 6 from the total length of 36 to find the tail. 36 - 6 = 30 so 30 is the length of the tail. :)

OpenStudy (anonymous):

30

OpenStudy (anonymous):

@JvBuenconcejo you cant just give the answer

OpenStudy (anonymous):

@NaeNae16 do you understand what I did?

OpenStudy (anonymous):

yes thank youu